Status Report

NASA Space Station On-Orbit Status 19 August 2005

By SpaceRef Editor
August 19, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 19 August 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  A light-duty day for the crew after yesterday’s spacewalk.

The crew’s sleep cycle remained shifted today, with wake-up 12:00 noon, but onset of sleep period returning to regular 5:30pm tonight.  Beginning tomorrow, the crew will be back on the normal schedule of 2:00am – 5:30pm (all times EDT).

Last night’s Orlan EVA-14 by was yet another success for the ISS Program.  

After opening DC1 airlock hatch #1 at 3:02pm EDT, CDR Krikalev and FE/SO Phillips spent 4h 58m outside the Russian segment (RS).  During this time, they – 

  • Removed & retrieved the “Biorisk” payload canister on the DC-1; 
  • Remove & retrieved MPAC & SEED panel #3 on the large diameter section (RO2) of the Service Module (SM); 
  • Installed a reserve television camera on the endcone of the SM Assembly Section to support ATV prox ops (the last ATV EVA/outfitting task);
  • Photographed & checked out the Kromka-3 contamination exposure experiment tablet attached to an SM handrail, leaving it in place (for lack of a containment bag); 
  • Swapped out the SKK-3-SM materials exposure container with the new SKK-5-SM; and 
  • Demated and retrieved the (deactivated and vented) Matryoshka experiment on SM RO. 

One task, removal of the Russian/US Strela 2 adapter from the EFGF (electrical flight grapple fixture) on the FGB module and its installation on the FRGF (flight releasable grapple fixture) on the PMA-3 (pressurized mating adapter #3), was scrubbed and deferred to a later EVA when the crew ran out of available Orlan time (the relocation of the end adapter to the PMA-3 will be required prior to Flight 12A.1 when the final set of SM meteoroid/orbital debris panels are delivered and stowed on PMA-3).  

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

The EVA ended with return to the DC1 and hatch closing at 8:00pm, for a total duration of 4h 58m.  This was the 62nd spacewalk in support of ISS assembly/maintenance and the 34th from the station itself (16 from “Pirs”, including one aborted).  With today’s EVA, ISS spacewalks have been staged by 42 NASA astronauts, 12 Russian cosmonauts, one Canadian, one Frenchman and one Japanese, running up a total of 373h 18m EVA time.  EVA-14 also was Phillips’ first and Krikalev’s eighth spacewalk (Previously, Sergei accumulated 36 hrs 10 min spacewalking experience on seven EVAs during his two back-to-back missions on the Russian MIR space station).

After return from the EVA and DC1 airlock repressurization from cabin air, the crew opened hatches and reentered the SM for ISS activation operations. Krikalev restored systems configurations in the DC1 and other RS modules to pre-EVA conditions, then installed the DC1 air ducts. 

Afterwards, John began reopening the USOS transfer hatches from the RS, then restored the ITCS (Internal Thermal Control System) and the OpsLAN network, transferring the PCS laptop back into the cabin. 

With post-EVA restoration complete, all ISS systems are back in a normal configuration, including comm links, OpsLAN computer network, IMV (inter-modular ventilation), TCCS (trace contaminant control subsystem), ITCS (internal thermal control system), and PCUs (plasma contactor units) back to Standby.  

A cabin atmosphere repress with N2 (nitrogen) from U.S. storage raised total cabin pressure to replace the pressure drop from the DC1 repress, and an O2 refresh from Progress 18 was also scheduled to maintain ppO2 (partial pressure oxygen).

Final close-out ops in the DC1 by Krikalev today included packing the retrieved containers of the SKK-3, MPAC & SEED and Biorisk-MSN payloads for return to Earth.

In addition, Sergei stowed the “Urolux” equipment used by the crew yesterday for the obligatory post-EVA session of the Russian crew health-monitoring program’s MO-9/Biochemical Urinalysis, and took the close-out readings of the “Pille-MKS” radiation dosimeters from the Orlan suits. 

As part of post-EVA close-out activities, FE Phillips deconfigured the Kodak DSC 760 EVA camera and stowed it with its elements back in the U.S. segment.

The CDR had ~2.5 hrs scheduled for the repair of the Russian Vozdukh carbon dioxide scrubber which failed last week (8/11).  The troubleshooting was to be essentially a swap-out of the two BVK vacuum valves, i.e., reversing the mounting positions (channel beds) of BVK-1 and BVK-2.  After remating the connectors, the system was to be activated in operational mode #5, supported by ground specialist tagup.

The FE meanwhile completed the regular daily SOZh/ECLSS maintenance plus ASU toilet system replacements, today also including the weekly inspection of the BRPK air/liquid condensate separator apparatus and the weekly routine checkup of the IP-1 airflow sensors in the various RS hatchways and FGB-to-Node tunnel. 

Afterwards, John conducted the weekly audit of CWCs (collapsible water containers), recording changes in CWC status over the last few days on a container cue card for subsequent downlink to MCC-H.

The U.S. CDRA (CO2 Removal Assembly) was successfully restarted after the EVA and is currently running nominally on one adsorbent/desiccant bed only (requiring periodic regeneration), until the stuck check valve in bed #2 re-seats so that CDRA can then be returned to its nominal Dual Bed configuration.   [Single-bed operations will continue until two cycles of proper check valve operation are observed.  CO2 levels are remaining well below Flight Rule (FR) limits. Presently CO2 is at 3.24 mmHg, the first FR “trigger” point for CO2 requiring action is 5.3 mmHg).]

No CEO (crew earth observations) photo targets uplinked today.

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 4:53am EDT [= epoch]):

  • Mean altitude — 353.7 km
  • Apogee height — 354.3 km
  • Perigee height — 353.0 km
  • Period — 91.61 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0000962
  • Solar Beta Angle — 31.6 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.72
  • Mean altitude loss in last 24 hours — 95 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 38562

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.